Please show me how to Write the equation in terms of y.
3X = 5 - y and 6x + 2y = 10

Found 2 solutions by stanbon, chitra:
Answer by stanbon(75887)

3X = 5 - y
Add y to both sides to get:
y+3x=5
Subtract 3x from both sides to get:
y=-3x+5
-------------------
and 6x + 2y = 10
Subtract 6x from both sides to get:
2y=-6x+10
Divide thru by 2 to get:
y=-3x+5
-----------
Cheers,
Stan H.

Answer by chitra(359)

The given equations are:

3x = 5 - y and 6x + 2y = 10
------------------------------------------------------------------------

3x = 5 - y

Adding y - 3x from both sides of equation (1), we get:

y = 5 - 3x This is the equation in terms of y.

-------------------------------------------------------------------------

6x + 2y = 10

subtracting 6x from both sides of the above equation, we get:

2y = 10 - 6x

Divide the entire equation by 2

y = 5 - 3x

This is the equation in terms of y